OpenStudy (anonymous):

A bag has 5 red marbles, 6 blue marbles and 4 black marbles. What is the probability of picking a blue marble, replacing it, and then picking a red marble?

OpenStudy (dumbcow):

6/15 * 5/15

OpenStudy (anonymous):

but these are the choices 8/75 3/5 2/15 4/45

OpenStudy (anonymous):

6/15*5/15 can be simplified to get to one of these answers : 6/15 = 2/5 (dividing both numerator and denominator by 3) 5/15 = 1/3 (dividing both numerator and denominator by 5) 2/5 * 1/3 will give you one of these answers

OpenStudy (anonymous):

2/15

OpenStudy (anonymous):

bingo

OpenStudy (dumbcow):

yeah you still have to simplify and reduce fractions...i was just showing you how to set it up probability = desired outcomes/total outcomes

OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a green marble, do not replace it, and pick a yellow marble? 1/15?

OpenStudy (anonymous):

close, but not quite. Can you show your working out so that I can explain what went wrong?

OpenStudy (anonymous):

oh my bad: (3 [yellow]/10 [total])×(4 [green]/9 [total minus one]) = 12/90 = 2/15

OpenStudy (anonymous):

That's exactly it

OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a green marble, do not replace it, and pick a red marble? 3/10

OpenStudy (anonymous):

Mathematically, what's different between this one and the last one?